# Investigating the Event-Shape Methods in Search for the Chiral Magnetic Effect in Relativistic Heavy Ion Collisions

### v1

@@ IV.3 Toy Model Verification
To test the hypothesis that resonance mixtures may influence the robustness of the ESS method, we use toy model simulations, where we can easily alter the event content to examine the consequences. In real data analysis, a fraction of $K_{S}$ hadrons decay into pions that are reconstructed as primordial particles, often defined experimentally as those with a reconstructed distance of closest approach (dca) from the reconstructed primary collision point within a certain cut, typically a few centimeters. In the STAR experiment, such a fraction is on the order of 2%, and is used in Toy Model I as default. To alter the event content, we vary the accepted fraction of $K_{S}$ to 10%, 20%, and 30%.
@@
In addition, we note in Fig. 2 that all the physics models contain a mass continuum of OS pairs in excess of SS pairs. We have therefore also included a mass continuum as described in Sect. III.6. The accepted $K_{S}$ fraction is kept as default. Adding the mass continuum is another way to alter the event content.
Figure A.12 and A.12 show the ESS and ESE results, respectively, from Toy Model I (default), Toy Model I (10% $K_{S}$), and Toy Model II (mass continuum).
Figure 7 summarizes the intercepts from the toy model studies. The intercepts from the ESS method appear to be different for the different event contents, verifying our hypothesis that the ESS method is sensitive to the event makeup because of the complex intertwining of the $\Delta\gamma$, $v_{2,{\rm single}}$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}$. The current statistics are too poor for the ESE method to draw conclusions, except for the default Toy Model I where the ESE intercept is consistent with zero.
@@
Figure 7: Toy model ESS and ESE intercepts. Shown are intercepts (left panel) and intercepts normalized by the inclusive $\Delta\gamma$ magnitude (right panel) from the ESS (blue) and ESE (red) methods with the toy models. Toy model inputs are from parameterizations to experimental data of 30–40% Au+Au collisions at 200 GeV. For ESS, the POI acceptance is $|\eta|<1$ and $20.2<p_{\rm T}<$ 2 GeV/$c$, and the $\hat{q}_{2,{\rm pair}}^{2}$ is always calculated from the same POIs. For ESE, the POI acceptance is $10.3<|\eta|<$ 1, and the $\hat{q}_{2}^{2}$ is calculated from particles in $|\eta|<0.3$, both with $20.2<p_{\rm T}<$ 2 GeV/$c$.
@@
The charge-dependent three-point correlator $\Delta\gamma$ is commonly used to search for the chiral magnetic effect (CME), a parity and charge-parity violating phenomenon predicted by quantum chromodynamics. The $\Delta\gamma$ correlator is contaminated by a major physics background induced by elliptic flow anisotropy $v_{2}$ in relativistic heavy ion collisions. The event-shape engineering (ESE) and event-shape selection (ESS) methods have been devised to search for the CME by projecting $\Delta\gamma$ to zero $v_{2}$ relying on dynamical fluctuations and statistical fluctuations of $v_{2}$, respectively. We have conducted a systematic study using four physical models (AVFD, AMPT, EPOS4, and HYDJET++) describing heavy ion collisions as well as a toy model. It is found, with no CME signal, that the ESS intercept ($\Delta\gamma_{\textsc{ess}}$) can be negative, zero, or positive depending on models and the ESE intercept ($\Delta\gamma_{\textsc{ese}}$) is mostly consistent with zero, as expected, albeit inherently large statistical uncertainties.
@@
It is further found that $\Delta\gamma_{\textsc{ess}}$ appears to depend on the details of the event content, for example, the relative abundances of resonances and the presence of a mass continuum. Given that the event-shape selection variable $\hat{q}_{2,{\rm pair}}^{2}$, the single particle $v_{2,{\rm single}}$ and the three-point correlator $\Delta\gamma$ in $\hat{q}_{2,{\rm pair}}^{2}$-classified events all depend on the particles of interest (POIs), which are made of not only primordial particles but also all decay products of electromagnetic- and strong-decay resonances and partial products of weak-decay resonances, we conjecture that the ESS intercept $\Delta\gamma_{\textsc{ess}}$ is a complicated function of all elements making up the event. These elements include resonance abundances relative to primordial pions, resonance $v_{2}$’s and $p_{\rm T}$ spectra, primordial particle $v_{2}$ and $p_{\rm T}$ spectra, etc. One may postulate a diagram such as in Fig. 8 where $\Delta\gamma_{\textsc{ess}}$ is depicted as a function of, for the sake of visualization, two elements of the event. For some combinations of the elements, $\Delta\gamma_{\textsc{ess}}$ is negative and for some others, $\Delta\gamma_{\textsc{ess}}$ is positive; for some lucky combinations, $\Delta\gamma_{\textsc{ess}}$ can happen to be zero. The issue is, unfortunately, that it is unknown, needless to say proven, what combinations of all those elements in the event would give the ideal $\Delta\gamma_{\textsc{ess}}=0$ in the absence of CME. It may be safe to say that Nature would unlikely give the ideal combinations to our collisions.
The ESE method, on the other side, comes with much more clarity–the intercept of flow-induced backgrounds $\Delta\gamma_{\textsc{ese}}$ must be zero–though it is statistically more demanding to achieve comparable precision. The statistics issue can easily be resolved, at least conceptually, by accumulating more data.
@@
Figure 8: Postulated behavior (for illustration only) of the ESS intercept $\Delta\gamma_{\textsc{ess}}$ on event elements of many dimensions, represented here for visualization purpose only by two, labeled as “resonance 1” and “resonance 2.” In the absence of CME, the value of $\Delta\gamma_{\textsc{ess}}$ can be negative or positive, and for some combinations of elements, can happen to be zero. This makes the ESS method a murky and highly model-dependent way for CME searches. This is because of the intertwining $\Delta\gamma$, $v_{2,{\rm single}}$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}$, all of which are calculated from the same POIs.
@@ Appendix A Simulation Plots
Figure A.1: AVFD ESS results. Three values of $n_{5}/s=0$ (left panel), 0.1 (center panel), and 0.2 (right panel) are implemented in the AVFD simulations of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV, with approximately $2\times 10^{7}$ events each for the 30–40% centrality. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<1$ and $20.2<p_{\rm T}<$ 2 GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10).
@@
Figure A.2: AVFD ESE results. Three values of $n_{5}/s=0$ (left panel), 0.1 (center panel), and 0.2 (right panel) are implemented in the AVFD simulations of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV, with approximately $2\times 10^{7}$ events each for the 30–40% centrality. The $\Delta\gamma$ is plotted as a function of $v_{2}$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $20.3<|\eta|<$ 2, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $20.2<p_{\rm T}<$ 2 GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10).
@@
Figure A.3: AMPT ESS results. Shown are three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV (upper panels) and at 27 GeV (lower panels) simulated by AMPT, with approximately $5\times 10^{6}$ events for each centrality at each energy. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<1$ and $20.2<p_{\rm T}<$ 2 GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10).
@@
Figure A.4: AMPT ESE results. Shown are three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV (upper panels) and at 27 GeV (lower panels) simulated by AMPT, with approximately $5\times 10^{6}$ events for each centrality at each energy. The $\Delta\gamma$ is plotted as a function of $v_{2}$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $10.3<|\eta|<$ 1, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $20.2<p_{\rm T}<$ 2 GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10).
@@
Figure A.5: EPOS ESS results. Shown are three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV (upper panels) and at 27 GeV (lower panels) simulated by EPOS4, with approximately $1.6\times 10^{6}$ and $8\times 10^{5}$ events for each centrality, respectively. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<2$ and $20.2<p_{\rm T}<$ 2 GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10).
@@
Figure A.6: EPOS ESE results. Shown are three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV (upper panels) and at 27 GeV (lower panels) simulated by EPOS4, with approximately $1.6\times 10^{6}$ and $8\times 10^{5}$ events for each centrality, respectively. The $\Delta\gamma$ is plotted as a function of $v_{2}$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $20.3<|\eta|<$ 2 and $20.2<p_{\rm T}<$ 2 GeV/$c$, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $20.2<p_{\rm T}<$ 2 GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10).
@@
Figure A.7: HYDJET++ ESS results. Shown are three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV (upper panels) and at 27 GeV (lower panels) simulated by HYDJET++, with approximately $4\times 10^{6}$ and $10^{7}$ events for each centrality, respectively. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<1$ and $20.2<p_{\rm T}<$ 2 GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10).
@@
Figure A.8: HYDJET++ ESE results. Shown are three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV (upper panels) and at 27 GeV (lower panels) simulated by HYDJET++, with approximately $4\times 10^{6}$ and $10^{7}$ events for each centrality, respectively. The $\Delta\gamma$ is plotted as a function of $v_{2}$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $10.3<|\eta|<$ 1 and $20.2<p_{\rm T}<$ 2 GeV/$c$, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $20.2<p_{\rm T}<$ 2 GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10).
Figure A.9: HYDJET++ ESE results (three-particle correlator). Shown are three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV (upper panels) and at 27 GeV (lower panels) simulated by HYDJET++, with approximately $4\times 10^{6}$ and $10^{7}$ events for each centrality, respectively. The $\Delta\gamma$ is plotted as a function of $v_{2}$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $10.3<|\eta|<$ 1, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $20.2<p_{\rm T}<$ 2 GeV/$c$. The three-particle correlator method is used to calculate $\Delta\gamma$ (Eq. 13) and the two-particle cumulant is used to calculate $\left\langle v_{2}\right\rangle$ (Eq. 8).
Figure A.10: HYDJET++ ESE results (forward/backward rapidity $\hat{q}_{2}^{2}$). Shown are three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV (upper panels) and at 27 GeV (lower panels) simulated by HYDJET++, with approximately $4\times 10^{6}$ and $10^{7}$ events for each centrality, respectively. The $\Delta\gamma$ is plotted as a function of $v_{2}$ in events binned in $\hat{q}_{2}^{2}\{2\}$. POIs are from acceptance $10.3<|\eta|<$ 1, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6) is computed from particles in forward/backward pseudorapidity regions of $43<|\eta|<$ 4, both with $20.2<p_{\rm T}<$ 2 GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ by Eqs. (2,3) and $\left\langle v_{2}\right\rangle$ by Eq. (10).

### v3

@@ IV.3 Toy Model Verification
One important difference among the models we studied is the various contributions from $K_{S}$. avfd and ampt keep the $K_{S}$ stable so there is no contamination from $K_{S}$ decay pions in the final state. epos4 and hydjet++, on the other hand, decay $K_{S}$ so there is maximum contamination in the final-state pions (as we have used all final-state pions of the model in our calculations). The intercepts from ampt at both 200 and 27 GeV are negative, and the intercept from avfd at 200 GeV is consistent with zero (avfd modeling at low energy is not available). The intercepts from epos4 at 200 and 27 GeV are both positive. The intercept from hydjet++ is consistent with zero at 200 GeV and positive at 27 GeV. In all models the strong-decay resonances (such as the $\rho^{0}$ resonance) are forced to decay. Relative particle abundances, for example, the $K_{S}$ over $\rho^{0}$ ratio, depends on energy; the lower the energy, the larger the low- to high-mass ratio. While very much dependent on models, it seems that $\Delta\gamma_{\textsc{ess}}$ increases with increasing $K_{S}/\rho^{0}$ ratio. It is thus conceivable, as resonances are a main background contribution to $\Delta\gamma$, that the relative resonance abundances could also be important for the ESS intercept.
@@
To test the hypothesis that resonance mixtures may influence the robustness of the ESS method, we use toy model simulations, where we can easily alter the event content to examine the consequences. In real data analysis, a fraction of $K_{S}$ hadrons decay into pions that are reconstructed as primordial particles, often defined experimentally as those with a reconstructed distance of closest approach (dca) from the reconstructed primary collision point to be within a certain cut, typically a few centimeters. In the STAR experiment, such a fraction is on the order of 2% [45], and is used as default in Toy Model I (described in Sect. III.5). To alter the event content, we vary the accepted fraction of $K_{S}$ to 10%, 20%, and 30%.
@@
In addition, we note in Fig. 2 that all the physics models contain a mass continuum of OS pairs in excess of SS pairs. We have therefore included a mass continuum in Toy Model II as described in Sect. III.6. The accepted $K_{S}$ fraction is kept at 2%. Adding the mass continuum is another way to alter the event content.
Figure A.14 shows the ESS results from Toy Model I (default 2% $K_{S}$), Toy Model I (10% $K_{S}$), and Toy Model II (mass continuum). It is interesting to notice that the dependence of $\Delta\gamma$ on $v_{2,{\rm single}}$ in those $\hat{q}_{2,{\rm pair}}^{2}$-binned events is not linear for Toy Model II. This suggests that the linear dependence hypothesis, which is essential for the linear extrapolation in ESS, may not be generally valid.
Figure A.14 shows the ESE results from Toy Model I (default 2% $K_{S}$), Toy Model I (10% $K_{S}$), and Toy Model II (mass continuum) with POIs from $10.3<|\eta|<$ 1 and $q_{2}^{2}$ from $|\eta|<0.3$. There seems to be an indication of nonlinear behavior in the Toy Model II result, similar to that in the ESS result in Fig. A.14 right panel.
@@
Figure 8: Toy model ESS and ESE intercepts. Shown are intercepts (upper panel) and intercepts divided by the inclusive $\Delta\gamma$ value (lower panel) from the ESS (blue points) and ESE (red and pink points) methods using the toy models. Toy model inputs are from parameterizations to experimental data of 30–40% Au+Au collisions at 200 GeV. For ESS, the POI acceptance is $|\eta|<1$ and $20.2<p_{\rm T}<$ 2 GeV/$c$, and the $\hat{q}_{2,{\rm pair}}^{2}$ is always calculated from the same POIs. For ESE, the POI acceptance is $10.3<|\eta|<$ 1, and the $\hat{q}_{2}^{2}$ is calculated from particles in $|\eta|<0.3$ (red points) or $43<|\eta|<$ 4 for three of the studied cases (pink points); The $p_{\rm T}$ range for both POIs and particles used for $\hat{q}_{2}^{2}$ calculations is $20.2<p_{\rm T}<$ 2 GeV/$c$. The rightmost red point is outside the frame.
Figure 8 summarizes the intercepts and the intercepts divided by the overall $\Delta\gamma$ values from the toy model studies. The intercepts from the ESS method (blue points) are all positive and appear to vary over the different event contents. This may confirm our hypothesis that the ESS method is sensitive to the details of the event makeup because of the complex intertwining of the $\Delta\gamma$, $v_{2,{\rm single}}$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}$.
The intercepts from the ESE method are shown in the red markers. The statistical precision is high because the toy models are computationally inexpensive. The ESE intercepts are consistent with zero for the two lower $K_{S}$-contamination cases studied in Toy Model I. However, the ESE intercepts from the two higher $K_{S}$-contamination cases of Toy Model I (positive) and the intercept from Toy Model II (negative) are inconsistent with zero. We postulate the cause to be nonflow correlations between POIs in $10.3<|\eta|<$ 1 and the $q_{2}^{2}$ calculated from $|\eta|<0.3$ in our toy model studies because resonance decay daughters can end up in both regions which are not far apart. (We have mentioned this nonflow effect in the $q_{2}^{2}$ event selection in our physics model discussion of Fig. 5, though the statistical uncertainties there are large.) In addition, the effect of adding a mass continuum is different from simply increasing the $K_{S}$ contamination, presumably because of the differing decay kinematics and opening angles.
To reduce this nonflow effect, we classify events in the ESE method using $q_{2}^{2}$ calculated from forward/backward pseudorapidity regions of $43<|\eta|<$ 4, far displaced from the POI’s region of $10.3<|\eta|<$ 1. The results are shown in Fig. A.14 for Toy Model I (20% $K_{S}$), Toy Model I (30% $K_{S}$), and Toy Model II (mass continuum). The ESE intercepts and the intercepts divided by the overall $\Delta\gamma$ values are shown in the pink markers in Fig. 8 with good statistical precision. These intercepts are consistent with zero, as one would expect in the absence of nonflow. These results highlight the importance to use $q_{2}^{2}$ calculated in pseudorapidity range far distanced from that of the POIs.
@@
The charge-dependent azimuthal correlator $\Delta\gamma$ is commonly used to search for the chiral magnetic effect (CME), a parity and charge-parity violating phenomenon predicted by quantum chromodynamics. The $\Delta\gamma$ correlator is contaminated by a major physics background induced by elliptic flow anisotropy $v_{2}$ in relativistic heavy ion collisions. The event-shape engineering (ESE) and event-shape selection (ESS) methods have been exploited to search for the CME by projecting $\Delta\gamma$ to the measured $v_{2}$ of vanishing value, relying on dynamical fluctuations and statistical fluctuations of $v_{2}$, respectively.
We have conducted a systematic study using four physics models (avfd, ampt, epos4, and hydjet++) describing heavy ion collisions. It is found, with no CME signal, that the ESS intercept ($\Delta\gamma_{\textsc{ess}}$) can be negative, zero, or positive depending on the models and multiplicities of the events, and the ESE intercept ($\Delta\gamma_{\textsc{ese}}$) is mostly consistent with zero, as expected, albeit inherently large statistical uncertainties.
@@
It is further found, using toy models, that $\Delta\gamma_{\textsc{ess}}$ appears to depend on the details of the event content, for example, the relative abundances of resonances and the presence of a mass continuum. Given that the event-shape selection variable $\hat{q}_{2,{\rm pair}}^{2}$, the single particle $v_{2,{\rm single}}$, and the three-point correlator $\Delta\gamma$ in $\hat{q}_{2,{\rm pair}}^{2}$-classified events all depend on the particles of interest (POIs), which are made of not only primordial particles but also all decay products of electromagnetic- and strong-decay resonances and partial products of weak-decay resonances, we conjecture that the ESS intercept $\Delta\gamma_{\textsc{ess}}$ is a complicated function of all elements that make up the event. These elements include resonance abundances relative to primordial pions, resonance $v_{2}$’s and $p_{\rm T}$ spectra, as well as primordial particle $v_{2}$ and $p_{\rm T}$ spectra. One may postulate a diagram such as in Fig. 9 where $\Delta\gamma_{\textsc{ess}}$ is depicted as a function of, for the sake of visualization, two elements of the event. For some combinations of the elements, $\Delta\gamma_{\textsc{ess}}$ is negative and for some others, $\Delta\gamma_{\textsc{ess}}$ is positive; for some accidental combinations, $\Delta\gamma_{\textsc{ess}}$ can happen to be zero. The issue is, unfortunately, that it is unknown, needless to say proven, what combinations of all those elements in the event would give the ideal $\Delta\gamma_{\textsc{ess}}=0$ in the absence of CME. This makes it impractical to search for CME using the ESS method.
@@
Figure 9: Postulated behavior (for illustration only) of the ESS intercept $\Delta\gamma_{\textsc{ess}}$ on event elements of many dimensions, represented here for visualization purpose only by two, labeled as “resonance 1” and “resonance 2.” In the absence of CME, the value of $\Delta\gamma_{\textsc{ess}}$ can be positive or negative, and for some combinations of elements, can happen to be zero. This is because of the intertwining $\Delta\gamma$, $v_{2,{\rm single}}$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}$, all of which are calculated from the same set of POIs. This makes the ESS method highly model-dependent and unpractical for rigorous CME searches.
The ESE method, on the other hand, has much more clarity: the $\Delta\gamma_{\textsc{ese}}$ intercept of flow-induced backgrounds must be zero in the absence of nonflow. However, the ESE method is not without pitfalls. There exist nonflow correlations between particles, and when those nonflow correlations are present between POIs and the $\hat{q}_{2}^{2}$ variable classifying events, then the ESE intercept can also be sensitive to non-CME physics. The issue is similar in nature to the ESS method–the relevant variables are now intertwined. One way to avoid/reduce such nonflow effects is to separate POIs and particles to calculate $\hat{q}_{2}^{2}$ well in pseudorapidity. Furthermore, the dynamical fluctuations in $v_{2}$ are typically not large, so a long extrapolation to the $v_{2}=0$ intercept is required in the ESE method. This results in a relatively large uncertainty in the $\Delta\gamma_{\textsc{ese}}$ intercept, and large statistics are required for a precision measurement by the ESE method. The statistics issue, however, can easily be resolved, at least conceptually, by accumulating more data.
@@ Appendix A Simulation Plots
Figure A.1: avfd ESS results. Three values of $n_{5}/s=0$ (left panel), 0.1 (center panel), and 0.2 (right panel) are implemented in the avfd simulations of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV, with approximately $2\times 10^{7}$ events each for the 30–40% centrality. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<1$ and $20.2<p_{\rm T}<$ 2 GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10). The red line is a first-order polynomial fit in the range of $0.20<v_{2,{\rm single}}<$ 0.2.
@@
Figure A.2: avfd ESE results. The same avfd data as in Fig. A.2 is used. The $\Delta\gamma$ is plotted as a function of $mean{v_{2}}$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $20.3<|\eta|<$ 2, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $20.2<p_{\rm T}<$ 2 GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10). The red line is a first-order polynomial fit to all data points.
@@
Figure A.3: ampt ESS results. Shown are three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV (upper panels) and at 27 GeV (lower panels) simulated by ampt, with approximately $4.2\times 10^{8}$ and $1.8\times 10^{9}$ events for each centrality at each energy. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<1$ and $20.2<p_{\rm T}<$ 2 GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10). The red line is a first-order polynomial fit in the range of $0.20<v_{2,{\rm single}}<$ 0.2
@@
Figure A.4: ampt ESE results. The same ampt data as in Fig. A.3 is used. The $\Delta\gamma$ is plotted as a function of $\left\langle v_{2}\right\rangle$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $10.3<|\eta|<$ 1, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $20.2<p_{\rm T}<$ 2 GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10). The red line is a first-order polynomial fit to all data points.
Figure A.5: ampt ESE results with forward/backward $\hat{q}_{2}^{2}\{2\}$. As same as Fig. A.5 except that the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in the forward/backward pseudorapidity region of $53<|\eta|<$ 5 for 200 GeV and $32<|\eta|<$ 3 for 27 GeV.
@@
Figure A.6: epos4 ESS results. Shown are three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV (upper panels) and at 27 GeV (lower panels) simulated by epos4, with approximately $1.6\times 10^{7}$ and $8.5\times 10^{6}$ events for each centrality, respectively. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<2$ and $20.2<p_{\rm T}<$ 2 GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10).
@@
Figure A.7: epos4 ESE results. The same epos4 data as in Fig. A.7 is used. The $\Delta\gamma$ is plotted as a function of $\left\langle v_{2}\right\rangle$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $20.3<|\eta|<$ 2, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $20.2<p_{\rm T}<$ 2 GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10).
@@
Figure A.8: hydjet++ ESS results. Shown are three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV (upper panels) and at 27 GeV (lower panels) simulated by hydjet++, with approximately $2.5\times 10^{8}$ and $5.7\times 10^{8}$ events for each centrality, respectively. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<1$ and $20.2<p_{\rm T}<$ 2 GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10).
@@
Figure A.9: hydjet++ ESE results. The same hydjet++ data as in Fig. A.9 is used. The $\Delta\gamma$ is plotted as a function of $\left\langle v_{2}\right\rangle$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $10.3<|\eta|<$ 1, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $20.2<p_{\rm T}<$ 2 GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10).
Figure A.10: hydjet++ ESE results (three-particle correlator). As same as Fig. A.9 except that the three-particle correlator method is used to calculate $\Delta\gamma$ (Eq. 13) and the two-particle cumulant is used to calculate $\left\langle v_{2}\right\rangle$ (Eq. 8), instead of using the known impact parameter direction in these calculations.
Figure A.11: hydjet++ ESE results (forward/backward rapidity $\hat{q}_{2}^{2}$). As same as Fig. A.9 except that the event selection variable $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6) is computed from particles in forward/backward pseudorapidity regions of $53<|\eta|<$ 5 for 200 GeV and $32<|\eta|<$ 3 for 27 GeV.